# Calibrating Reasoning in Language Models with Internal Consistency

### v1

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in various reasoning tasks, aided by techniques like chain-of-thought (CoT) prompting that elicits verbalized reasoning. However, LLMs often generate text with obvious mistakes and contradictions, raising doubts about their ability to robustly process and utilize generated rationales. In this work, we investigate CoT reasoning in LLMs through the lens of internal representations, focusing on how these representations are influenced by generated rationales. Our preliminary analysis reveals that while generated rationales improve answer accuracy, inconsistencies emerge between the model’s internal representations in middle layers and those in final layers, potentially undermining the reliability of their reasoning processes. To address this, we propose internal consistency as a measure of the model’s confidence by examining the agreement of latent predictions decoded from intermediate layers. Extensive empirical studies across different models and datasets demonstrate that internal consistency effectively distinguishes between correct and incorrect reasoning paths. Motivated by this, we propose a new approach to calibrate CoT reasoning by up-weighting reasoning paths with high internal consistency, resulting in a significant boost in reasoning performance. Further analysis uncovers distinct patterns in attention and feed-forward modules across layers, providing insights into the emergence of internal inconsistency. In summary, our results demonstrate the potential of using internal representations for self-evaluation of LLMs.
@@ 1 Introduction
Figure 1: An illustration of internal consistency. Given a true-or-false question with the ground truth being true, we elicit latent predictions (i.e., predictions decoded from intermediate layers, represented by “T” for true and “F” for false) from the answer token of CoT reasoning paths. By defining internal consistency as the agreement of latent predictions with the final stated one, we observe a high correlation between internal consistency and prediction accuracy, which aids in calibrating CoT reasoning. Note that the figure on the right is synthesized for illustration purposes, but the distributions reflect actual observations as shown in Figure 3.
@@
Based on this observation, we propose internal consistency (illustrated in Figure 1) as a measure of the model’s confidence in the reasoning process. Inspired by previous works on eliciting latent predictions from intermediate layers (nostalgebraist, 2020; Belrose et al., 2023), we examine internal consistency by assessing the agreement of latent predictions decoded from intermediate layers. Unlike methods that require additional training and human annotations (Ye and Durrett, 2022; Khalifa et al., 2023), internal consistency provides a reliable and off-the-shelf self-evaluation for CoT reasoning.
@@
This work makes three main contributions. 1) We identify the emergence of inconsistencies between intermediate and final layer representations in LLM reasoning, highlighting a potential issue where CoT reasoning leads to higher uncertainty. 2) We propose internal consistency as a novel measure to evaluate the model’s confidence and calibrate CoT reasoning by up-weighting paths with high internal consistency. 3) We provide insights into the cause of internal inconsistency in CoT reasoning by analyzing Transformer components (i.e., attention and feed-forward networks) across layers. We believe our results show promise in leveraging internal representations to enhance reasoning in LLMs.
@@ 3.2 Calibration Measurement with Internal Consistency
$\operatorname{InternalConsistency}(\mathbf{x},\hat{\mathbf{y}})=\frac{1}{L-1}\sum_{\ell=1}^{L-1}\mathbbm{1}\{\hat{\mathbf{y}}^{\ell}=\hat{\mathbf{y}}^{L}\}.$
@@ 4.1 Setup
Figure 3: Internal consistency is a reliable measure of prediction confidence in CoT reasoning. The four panels display: 1) the effect of different prompting techniques on the model’s internal consistency; 2) the distribution discrepancy of internal consistency between correct and incorrect model predictions; 3) pattern variations in agreement values (representing the ratio of data instances where the latent predictions match the final predictions) across layers; and 4) a calibration plot with bins according to the model’s internal consistency on the x-axis and the accuracy within each bin on the y-axis. Results are averaged over all models and datasets. See Appendix C for the full results.
@@ 4.2 Internal Consistency is a Good Calibration Measure
Our analysis begins by examining patterns of internal consistency across different layers of LLMs. As illustrated in Figure 6, latent predictions exhibit variable convergence during inference; notably, there are significant increases in consistency in the middle and final layers, while the early layers are characterized by considerable noise. Additionally, these patterns vary across models and datasets. Smaller models, such as Llama-2-7B, show high variability in latent predictions across layers, indicating less stability in their reasoning processes. In contrast, larger models like Mixtral 8$\times$7B demonstrate more robust internal reasoning, as evidenced by more reliable convergence of latent predictions. For BoolQ, the inconsistency is less pronounced compared to other datasets that require more complex symbolic and logical reasoning.
@@
4.3 Internal Consistency Improves Chain-of-Thought Reasoning
Building on observed correlations between internal consistency and prediction accuracy, we further explore the use of internal consistency to enhance reasoning performance. Our focus is on CoT prompting, which generally decreases internal consistency as shown in Section 4.2. This decrease is possibly attributable to the generated reasoning paths that incorporate incorrect rationales, which confuse the model to make inconsistent predictions across layers. Given these insights, we aim to investigate whether we could leverage internal consistency to boost CoT reasoning.
We build on the self-consistency (SC) approach (Wang et al., 2022), which utilizes an ensemble of multiple sampled reasoning paths to increase accuracy. Our proposed method, which we term SC+IC, integrates internal consistency to weight these reasoning paths accordingly. Specifically, if a reasoning path demonstrates high internal consistency in its answer prediction, the probability of accepting its corresponding answer as the final answer increases. We accumulate the sum of internal consistency scores across all reasoning paths for each answer and choose the answer with the highest sum. For baselines, we also compare with vanilla CoT and another logit-based approach SC+$\Delta$ (Wang and Zhou, 2024) that selects the final answer based on a confidence measure $\Delta_{k}=p\left(\hat{\mathbf{y}}^{1}\mid\mathbf{x}\right)-p\left(\hat{\mathbf{y}}^{2}\mid\mathbf{x}\right)$, where $\hat{\mathbf{y}}^{1}$ and $\hat{\mathbf{y}}^{2}$ represent the top two tokens for answer prediction of the $k$-th path.
@@ 4.3 Internal Consistency Improves Chain-of-Thought Reasoning
Table 1: Internal consistency boosts CoT reasoning. We report calibrated accuracy and standard deviation averaged across runs of 10 different random seeds. For SC and its variants, results are obtained with 40 sampled reasoning paths.
Method BoolQ CoinFlip PrOntoQA ProofWriter Mean
Llama-2-7B CoT 67.1 $\pm$ 0.66 71.2 $\pm$ 1.80 51.7 $\pm$ 1.68 52.4 $\pm$ 1.17 60.6
SC 70.0 $\pm$ 0.05 76.0 $\pm$ 0.10 49.0 $\pm$ 0.33 52.2 $\pm$ 0.10 61.8
SC+$\Delta$ 69.9 $\pm$ 0.00 76.2 $\pm$ 0.00 50.2 $\pm$ 0.00 52.4 $\pm$ 0.00 62.2
SC+IC 70.7 $\pm$ 0.00 77.6 $\pm$ 0.09 49.0 $\pm$ 0.00 52.2 $\pm$ 0.00 62.4
Llama-2-13B CoT 78.5 $\pm$ 0.53 78.9 $\pm$ 1.02 53.6 $\pm$ 1.72 60.8 $\pm$ 1.59 67.9
SC 80.8 $\pm$ 0.12 81.0 $\pm$ 0.23 52.6 $\pm$ 0.35 61.0 $\pm$ 0.29 68.9
SC+$\Delta$ 81.2 $\pm$ 0.00 81.8 $\pm$ 0.00 55.0 $\pm$ 0.00 61.6 $\pm$ 0.00 69.9
SC+IC 81.4 $\pm$ 0.02 84.0 $\pm$ 0.00 55.4 $\pm$ 0.00 61.8 $\pm$ 0.00 70.6
Mistral-7B CoT 74.0 $\pm$ 0.49 89.8 $\pm$ 1.36 55.3 $\pm$ 1.61 65.6 $\pm$ 0.91 71.2
SC 76.8 $\pm$ 0.09 95.2 $\pm$ 0.22 58.0 $\pm$ 0.35 68.6 $\pm$ 0.27 74.6
SC+$\Delta$ 77.9 $\pm$ 0.00 95.2 $\pm$ 0.00 59.6 $\pm$ 0.00 66.8 $\pm$ 0.00 74.9
SC+IC 78.1 $\pm$ 0.00 96.4 $\pm$ 0.00 59.8 $\pm$ 0.00 70.0 $\pm$ 0.00 76.1
Mixtral-8$\times$7B CoT 78.8 $\pm$ 0.42 98.8 $\pm$ 0.35 57.4 $\pm$ 1.39 72.3 $\pm$ 1.69 76.8
SC 81.6 $\pm$ 0.07 99.4 $\pm$ 0.00 61.6 $\pm$ 0.31 75.6 $\pm$ 0.30 79.6
SC+$\Delta$ 82.6 $\pm$ 0.00 99.4 $\pm$ 0.00 63.4 $\pm$ 0.00 75.0 $\pm$ 0.00 80.1
SC+IC 81.7 $\pm$ 0.00 99.4 $\pm$ 0.00 63.4 $\pm$ 0.00 78.0 $\pm$ 0.00 80.6
Figure 4: Internal consistency brings more performance gain in tasks involving complex reasoning. The figure shows calibrated accuracy as a function of the number of votes for three types of tasks: reading comprehension (BoolQ), symbolic reasoning (CoinFlip), and logical reasoning (PrOntoQA and ProofWriter). Results are averaged over all models and random seeds.
@@ 4.4 Investigating the Role of Transformer Components in Internal Inconsistency
We analyze the average attention weights across all attention heads at each layer, focusing on the answer tokens. Consistent with the CoT process, we segment the processed tokens into three parts: the context, the query, and the generated rationale. Further details on the methodology and calculations can be found in Appendix B.4.
@@
For FFN layers, we use a probe vector trained on the model’s last hidden state of the answer token with respect to the model’s output. We compute cosine similarities between this probe vector and the value vectors (Geva et al., 2022; Lee et al., 2024), each corresponding to a column in the second FFN matrix. Our analysis focuses on the vectors with the top 0.1% of cosine similarity values, which play a pivotal role in forming final predictions. For more details, please see Appendix B.5.
@@ 5 Related Work
Although the reasoning capabilities of LLMs have been greatly enhanced with techniques like CoT reasoning (Wei et al., 2022; Yao et al., 2022; Merrill and Sabharwal, 2024), previous investigations into the faithfulness of CoT reasoning have shown that the model’s generated rationales often are not consistent with their predictions (Lyu et al., 2023; Lanham et al., 2023). These studies primarily focus on the alignment between a model’s verbalized reasoning and its outcomes, without delving into the model’s internal reasoning processes. In contrast, our work provides new insights into how unfaithfulness emerges internally during CoT reasoning and proposes solutions to calibrate reasoning with internal consistency.
@@ B.2 Main Experiments
Decoding

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in various reasoning tasks, aided by techniques like chain-of-thought prompting that elicits verbalized reasoning. However, LLMs often generate text with obvious mistakes and contradictions, raising doubts about their ability to robustly process and utilize generated rationales. In this work, we investigate reasoning in LLMs through the lens of internal representations, focusing on how these representations are influenced by generated rationales. Our preliminary analysis reveals that while generated rationales improve answer accuracy, inconsistencies emerge between the model’s internal representations in middle layers and those in final layers, potentially undermining the reliability of their reasoning processes. To address this, we propose internal consistency as a measure of the model’s confidence by examining the agreement of latent predictions decoded from intermediate layers. Extensive empirical studies across different models and datasets demonstrate that internal consistency effectively distinguishes between correct and incorrect reasoning paths. Motivated by this, we propose a new approach to calibrate reasoning by up-weighting reasoning paths with high internal consistency, resulting in a significant boost in reasoning performance. Further analysis uncovers distinct patterns in attention and feed-forward modules across layers, providing insights into the emergence of internal inconsistency. In summary, our results demonstrate the potential of using internal representations for self-evaluation of LLMs. Our code is available at github.com/zhxieml/internal-consistency.
@@ 1 Introduction
Figure 1: An illustration of internal consistency. Given a true-or-false question with the ground truth being true, we elicit latent predictions (i.e., predictions decoded from intermediate layers, represented by “T” for true and “F” for false) from the answer token of reasoning paths. By defining internal consistency as the agreement of latent predictions with the final stated one, we observe a high correlation between internal consistency and prediction accuracy, which aids in calibrating reasoning. Note that the figure on the right is synthesized for illustration purposes, but the distributions reflect actual observations as shown in Figure 3.
@@
Based on this observation, we propose internal consistency (illustrated in Figure 1) as a measure of the model’s confidence in the reasoning process. Inspired by previous works on eliciting latent predictions from intermediate layers (nostalgebraist, 2020; Belrose et al., 2023), we examine internal consistency by assessing the agreement of latent predictions decoded from intermediate layers. Unlike methods that require additional training and human annotations (Ye and Durrett, 2022; Khalifa et al., 2023), internal consistency provides a reliable and off-the-shelf self-evaluation for reasoning.
@@
This work makes three main contributions. 1) We identify the emergence of inconsistencies between intermediate and final layer representations in LLM reasoning, highlighting a potential issue where CoT reasoning leads to higher uncertainty. 2) We propose internal consistency as a novel measure to evaluate the model’s confidence and calibrate reasoning by up-weighting paths with high internal consistency. 3) We provide insights into the cause of internal inconsistency in reasoning by analyzing Transformer components (i.e., attention and feed-forward networks) across layers. We believe our results show promise in leveraging internal representations to enhance reasoning in LLMs.
@@ 3.2 Calibration Measurement with Internal Consistency
$\operatorname{InternalConsistency}(\mathbf{x},\hat{\mathbf{y}})=\frac{1}{L-1}\sum_{\ell=1}^{L-1}\mathbbm{1}\{\hat{\mathbf{y}}^{\ell}=\hat{\mathbf{y}}^{L}\}.$ (2)
@@ 4.1 Setup
Figure 3: Internal consistency is a reliable measure of prediction confidence in CoT reasoning. From left to right: 1) the effect of different prompting techniques on the model’s internal consistency; 2) the distribution discrepancy of internal consistency between correct and incorrect model predictions; 3) pattern variations in agreement values (representing the ratio of data instances where the latent predictions match the final predictions) across layers; and 4) a calibration plot with bins according to the model’s internal consistency on the x-axis and the accuracy within each bin on the y-axis. Results are averaged over all models and datasets. See Appendix C for the full results.
@@ 4.2 Internal Consistency is a Good Calibration Measure
Our analysis begins by examining patterns of internal consistency across different layers of LLMs. As illustrated in Figure 6 (Appendix C), latent predictions exhibit variable convergence during inference; notably, there are significant increases in consistency in the middle and final layers, while the early layers are characterized by considerable noise. Additionally, these patterns vary across models and datasets. Smaller models, such as Llama-2-7B, show high variability in latent predictions across layers, indicating less stability in their reasoning processes. In contrast, larger models like Mixtral 8$\times$7B demonstrate more robust internal reasoning, as evidenced by more reliable convergence of latent predictions. For BoolQ, the inconsistency is less pronounced compared to other datasets that require more complex symbolic and logical reasoning.
@@
4.3 Enhancing Reasoning with Internal Consistency
Our analysis reveals a strong correlation between internal consistency and prediction accuracy, suggesting its potential as a mechanism for improving reasoning performance. As demonstrated in Section 4.2, CoT prompting, while effective, tends to decrease internal consistency. This decrease is possibly attributable to the generated reasoning paths that incorporate incorrect rationales, which confuse the model to make inconsistent predictions across layers. To investigate the generality of this phenomenon and its potential solutions, we also least-to-most (L2M) prompting (Zhou et al., 2022), which decomposes complex problems into simpler sub-problems for reasoning. For a detailed description of our least-to-most decomposition approach, please refer to Appendix B.3.
To incorporate internal consistency, we build on the self-consistency (SC) approach (Wang et al., 2022), which utilizes an ensemble of multiple sampled reasoning paths to increase accuracy. Our proposed method, which we term SC+IC, integrates internal consistency to weight these reasoning paths accordingly. Specifically, if a reasoning path demonstrates high internal consistency in its answer prediction, the probability of accepting its corresponding answer as the final answer increases. We accumulate the sum of internal consistency scores across all reasoning paths for each answer and choose the answer with the highest sum. For baselines, we also compare with greedy decoding Greedy and a logit-based approach SC+$\Delta$ (Wang and Zhou, 2024) that selects the final answer based on a confidence measure $\Delta_{k}=p\left(\hat{\mathbf{y}}^{1}\mid\mathbf{x}\right)-p\left(\hat{\mathbf{y}}^{2}\mid\mathbf{x}\right)$, where $\hat{\mathbf{y}}^{1}$ and $\hat{\mathbf{y}}^{2}$ represent the top two tokens for answer prediction of the $k$-th path.
@@ 4.3 Enhancing Reasoning with Internal Consistency
Table 1: Internal consistency improves reasoning performance across diverse tasks. We report calibrated accuracy averaged across runs of 10 different random seeds. For self-consistency (SC) and its variants, results are obtained with 40 sampled reasoning paths.
BoolQ CoinFlip PrOntoQA ProofWriter Mean
CoT L2M CoT L2M CoT L2M CoT L2M CoT L2M
Llama-2-7B
Greedy 67.1 61.4 71.2 81.9 51.7 51.1 52.4 60.7 60.6 63.8
SC 70.0 73.2 76.0 89.2 49.0 50.0 52.2 66.8 61.8 69.8
SC+$\Delta$ 69.9 73.0 76.2 90.1 50.2 50.1 52.4 68.2 62.2 70.3
SC+IC 70.7 73.4 77.6 90.2 49.0 51.2 52.2 69.0 62.4 71.0
SC+IC (tune) 71.5 73.5 78.6 93.9 50.8 55.7 54.1 70.7 63.8 73.4
SC+IC (transfer) 71.5 73.5 78.7 93.9 50.8 55.7 53.9 70.7 63.8 73.4
Llama-2-13B
Greedy 78.5 71.3 78.9 90.5 53.6 54.0 60.8 78.1 67.9 73.5
SC 80.8 78.3 81.0 94.2 52.6 55.4 61.0 88.6 68.9 79.1
SC+$\Delta$ 81.2 78.9 81.8 94.2 55.0 57.6 61.6 87.9 69.9 79.6
SC+IC 81.4 78.7 84.0 93.2 55.4 57.0 61.8 88.6 70.6 79.4
SC+IC (tune) 81.4 80.4 86.2 96.5 54.5 56.6 62.9 89.6 71.2 80.8
SC+IC (transfer) 81.4 80.2 86.4 96.3 54.4 56.2 62.8 89.7 71.2 80.6
Mistral-7B
Greedy 74.0 72.7 89.8 99.1 55.3 53.4 65.6 73.3 71.2 74.6
SC 76.8 78.9 95.2 99.4 58.0 51.4 68.6 84.0 74.6 78.4
SC+$\Delta$ 77.9 79.5 95.2 99.4 59.6 52.7 66.8 85.3 74.9 79.2
SC+IC 78.1 78.9 96.4 99.4 59.8 53.2 70.0 85.0 76.1 79.1
SC+IC (tune) 77.8 79.5 97.7 99.9 60.4 56.6 71.2 88.6 76.8 81.2
SC+IC (transfer) 77.8 79.4 97.7 99.9 60.5 56.5 71.1 88.7 76.8 81.1
Mixtral-8$\times$7B
Greedy 78.8 78.2 98.8 99.4 57.4 57.9 72.3 79.4 76.8 78.7
SC 81.6 85.3 99.4 100.0 61.6 56.8 75.6 90.2 79.6 83.1
SC+$\Delta$ 82.6 85.6 99.4 100.0 63.4 57.9 75.0 90.6 80.1 83.6
SC+IC 81.7 85.6 99.4 100.0 63.4 56.4 78.0 90.6 80.6 83.2
SC+IC (tune) 81.8 86.1 99.9 100.0 63.8 59.3 78.8 92.1 81.1 84.4
SC+IC (transfer) 81.8 86.1 100.0 100.0 63.8 59.5 78.9 92.1 81.1 84.4
4.4 Layer-weighted Aggregation Finds Transferable Patterns
While results of SC+IC have demonstrated the effectiveness of leveraging latent predictions for reasoning, the relative importance of different layers remains unexplored. Recent studies have shown that certain intermediate layers specialize in specific types of reasoning (Yang et al., 2024). These observations suggest that treating all layers equally when computing internal consistency might be suboptimal. Motivated by these insights, we introduce an aggregation strategy for internal consistency with two variants: SC+IC (tune) and SC+IC (transfer). Specifically, instead of assigning equal weights to each layer in Equation 2, we consider a learned weight vector $\mathbf{w}\in\mathbb{R}^{L}$ (where $L$ is the number of layers) for aggregation. For SC+IC (tune), we optimize layer weights using 500 held-out samples per dataset. The weights are learned using Adam optimizer with a learning rate of 0.01 for 1,000 iterations. SC+IC (transfer) evaluates cross-task generalization by applying weights tuned on PrOntoQA to other datasets.
The results are presented in Table 1 and Figure 4. Despite introducing only $L$ parameters, this weighted approach yields substantial improvements. Notably, the strong performance of SC+IC (transfer) across diverse tasks indicates that optimal layer importance patterns generalize well, demonstrating the broad applicability of internal consistency for reasoning tasks.
Figure 4: Internal consistency brings larger gains for complex reasoning tasks. The figure shows calibrated accuracy as a function of the number of votes for three types of tasks: reading comprehension (BoolQ), symbolic reasoning (CoinFlip), and logical reasoning (PrOntoQA and ProofWriter). Results are averaged over all models and random seeds.
@@ 4.5 Investigating the Role of Transformer Components in Internal Inconsistency
We analyze the average attention weights across all attention heads at each layer, focusing on what the answer tokens attend to. Consistent with the CoT process, we segment the processed tokens into three parts: the context, the query, and the generated rationale. Further details on the methodology and calculations can be found in Appendix B.5.
@@
For FFN layers, we use a probe vector trained on the model’s last hidden state of the answer token with respect to the model’s output. We compute cosine similarities between this probe vector and the value vectors (Geva et al., 2022; Lee et al., 2024), each corresponding to a column in the last FFN matrix of the layer. Our analysis focuses on the vectors with the top 0.1% of cosine similarity values, which play a pivotal role in forming final predictions. For more details, please see Appendix B.6.
@@ 5 Related Work
Although the reasoning capabilities of LLMs have been greatly enhanced with techniques like CoT reasoning (Wei et al., 2022; Yao et al., 2022; Xia et al., 2024), previous investigations into the faithfulness of CoT reasoning have shown that the model’s generated rationales often are not consistent with their predictions (Lyu et al., 2023; Lanham et al., 2023). These studies primarily focus on the alignment between a model’s verbalized reasoning and its outcomes, without delving into the model’s internal reasoning processes. In contrast, our work provides new insights into how unfaithfulness emerges internally during CoT reasoning and proposes solutions to calibrate reasoning with internal consistency.
@@ B.2 Main Experiments
B.3 Least-to-Most Prompting
In addition to chain-of-thought prompting, we evaluated our approach using least-to-most prompting Zhou et al. [2022] across four datasets. Due to the absence of existing least-to-most prompting templates for these datasets, we developed a systematic process to generate and validate appropriate prompts.
For each dataset, we selected 4 representative instances as our prompt engineering samples. We utilized GPT-4o to generate initial least-to-most decompositions for these instances, followed by manual inspection and refinement. Table 4 presents examples from each dataset demonstrating our least-to-most decomposition approach.
Table 4: Examples of least-to-most prompting decomposition across different datasets.
Dataset Context Decomposition Steps
BoolQ System of a Down, sometimes shortened to System and abbreviated as SOAD, is an Armenian-American heavy metal band from Glendale, California, formed in 1994. The band currently consists of Serj Tankian (lead vocals, keyboards), Daron Malakian (vocals, guitar), Shavo Odadjian (bass, backing vocals) and John Dolmayan (drums). 1. Who are the members of System of a Down? 2. What are the roles of each member? 3. How many members are listed as vocalists?
CoinFlip A coin is heads up. Ka flips the coin. Sherrie flips the coin. 1. What is the starting position of the coin? 2. What happens to the coin after each flip?
PrOntoQA Every tumpus is sour. Tumpuses are rompuses. Every rompus is small. Every rompus is an impus. Impuses are floral. Impuses are dumpuses. Dumpuses are not kind. Every dumpus is a yumpus. Each yumpus is feisty. Yumpuses are numpuses. Numpuses are not opaque. Each numpus is a zumpus. Every wumpus is opaque. Every zumpus is temperate. Each zumpus is a jompus. Every jompus is dull. Jompuses are vumpuses. Alex is a dumpus. 1. What are the properties of a dumpus? 2. Is Alex a dumpus? 3. Is a dumpus opaque?
ProofWriter Fiona is blue. Harry is cold. Harry is white. If Harry is blue and Harry is green then Harry is not round. If something is green then it is young. All white things are young. If something is green and not white then it is not blue. Young, round things are not furry. If something is white and young then it is round. If something is young and not cold then it is round. If something is green and not young then it is not round. 1. What are Harry’s characteristics? 2. What does being white imply about being young? 3. What does being white and young imply about being round?
Our decomposition strategy focuses on breaking down complex reasoning tasks into simpler, sequential steps. This systematic approach helps the model handle complex reasoning tasks by addressing simpler sub-problems sequentially.
Generation